How do I find a side length of a polygon if I only know the area? I specifically need to find the side lengths of a regular pentagon with an area of 49^2 meters.
Answer by Alan3354(69443)

How do I find a side length of a polygon if I only know the area? I specifically need to find the side lengths of a regular pentagon with an area of 49^2 meters.
-----------------
If you mean 49 sq meters:
---
For any regular polygon:
Area = ns^2*cot(180/n)/4 where n = # of sides and s = side length
49 = 5s^2*cot(36)/4
s^2 = (196/5)*tan(36)
s =~ 5.3367 meters
=====================
w/o using the formula:
---
The pentagon is 5 isosceles triangles, each with a central angle of 72 degs and an area of 49/5 sq meters.